OpenStudy (anonymous):

Confused, Algebra 2? Question 1: What I did - STEP 1. 5a + 5b; a = -6 and b = -5 Plug in for the given variables a (-6) and b (-5) (5 * -6) + (5 * -5) STEP 2. (5 • -6) + (5 • -5) Simplify using PEMDAS (when subtracting a negative number from another negative numnber you add the amount being subtracted as a positive number. -30 + 25 STEP 3. -30 + 25 Solve. -5 What My Teacher Said - She gave me partial credit and her comment was "-30 - 25 = ?"

OpenStudy (bibby):

That's because you added 25. (5 • -6) + (5 • -5) = (-30) + (-25) -30 - 25 = -55

OpenStudy (amistre64):

when 'adding' a negative number to another negative number ... think of a negative as a hole, and you are making the hole deeper. -a + -b = -a-b = -(a+b)